Selected financial data of two competitors, Target and Wal-Mart, are presented here. (All dollars are in millions.) Suppose the data were taken from the 2022 financial statements of each company. 

                            Target        Wal-Mart

                           (1/31/22)     (1/31/22)

                            Income Statement Data for

                                   Year

 Net sales                    $65,000    $421,000

 Cost of goods sold            45,500     319,960

 Selling and administrative expenses        16,250        88,410

 Interest expense                650        4,210

 Other income                     20        4,300

 Income taxes                    917        4,452

 Net income                    $1,703      $8,268


                                                   Balance Sheet Data (End of

                                                           Year)

 Current assets                                     $19,680         $46,980

 Noncurrent assets                                   33,000         122,000

 Total assets                                       $52,680        $168,980

 Current liabilities                                $12,000         $ 54,000

 Long-term liabilities                               24,876          47,388

 Total stockholders' equity                          15,804          67,592

 Total liabilities and stockholders' equity         $52,680        $168,980

 Net cash provided by operating activities           $4,600         $23,000

 Cash paid for capital expenditures                  $3,400         $11,700

 Dividends declared and paid on common stock          $530           $3,500

 Weighted-average common shares outstanding (millions) 650            2,650


For each company, compute these values and ratios. (All dollars are in millions.) (Round Current ratio and Earnings per share to 2 decimal places, e.g. 15.25 and Debt to assets ratio to 1 decimal place, e.g. 78.9%. If answer is negative enter it with a negative sign preceding the number e.g.-15,000 or in parentheses e.g. (15,000).) 


(a) Working capital 

(b) Current ratio 

(c) Debt to assets ratio 

(d) Free cash flow 

(e) Earnings per share 

(f) Which company has better liquidity? 

Which company has better solvency?
Answer #1
Target Wal-mart
Working capital = Current assets - Current liabilities $19,680-12,000 = $7,680 $46,980-54,000 = -$7,020
Current ratio = Current assets/Current liabilities $19,680/12,000 = 1.64 $46,980/54,000 = 0.87
Debt to assets ratio = Total liabilities/Total assets $36,876/52,680 = 0.7 $101,388/168,980 = 0.6
Free cash flow = Net cash provided by operating activities-Cash paid for capital expenditure $4,600-3,400 = $1,200 $23,000-11,700 = $11,300
Earnings per share = Net income /Weighted average common shares outstanding $1,703/650 = $2.62 $8,268/2,650 = $3.12

f.

Which company has better liquidity - Target

Which company has better solvency - Target